### v1

@@ Appendix F Interpretability
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Medial Belt Complex Area 7PC $0.002$ HC 3.728
Area TA2 (temporal region A, area 2) Area 7PC $0.005$ HC 3.157
ParaBelt Complex Area Frontal Opercular 5 $0.005$ HC 3.119
Dorsal area 6 Area Frontal Opercular 5 $0.009$ HC 2.781
ParaHippocampal Area 1 Area 7PC $0.009$ HC 2.752
Auditory 4 Complex Area 7PC $0.011$ HC 2.739
ParaHippocampal Area 1 Area posterior 9-46 ventral $0.010$ HC 2.729
Auditory 4 Complex Area Frontal Opercular 5 $0.011$ HC 2.696
Area PH Area Frontal Opercular 5 $0.004$ HC 2.570
Superior 6-8 Transitional Area Area 7PC $0.013$ HC 2.551
Posterior Insular Area 2 Area Frontal Opercular 5 $0.014$ HC 2.546
Medial Belt Complex Area Lateral IntraParietal dorsal $0.005$ HC 2.541
ParaHippocampal Area 1 Area IFJa $0.014$ HC 2.520
Medial Belt Complex Area posterior 9-46 ventral $0.014$ HC 2.518
Area anterior 9-46 ventral Area 7PC $0.014$ HC 2.515
Area OP2-3-VS (Ventral-Superior part of 2-3 area of Parietal Operculum) Area Frontal Opercular 5 $0.014$ HC 2.503
Medial Belt Complex Area IFJa $0.017$ HC 2.444
Area 8B Lateral Area 7PC $0.006$ HC 2.440
Area 6 medial anterior Area Lateral IntraParietal dorsal $0.005$ HC 2.436
ParaBelt Complex Area 7PC $0.017$ HC 2.433
Middle Temporal Area Area 7PC $0.017$ HC 2.402
Area 52 Area Lateral IntraParietal dorsal $0.006$ HC 2.398
Area 46 Area Frontal Opercular 5 $0.018$ HC 2.394
Medial Belt Complex Frontal Opercular Area 2 $0.018$ HC 2.379
Dorsal area 6 Area Lateral IntraParietal dorsal $0.006$ HC 2.374
Area 6 medial posterior Area Lateral IntraParietal dorsal $0.019$ HC 2.364
Posterior Insular Area 2 Area 7PC $0.019$ HC 2.357
Area OP4-PV (Posterior-Ventral part of area 4 of Parietal Operculum) Area Frontal Opercular 5 $0.020$ HC 2.335
Inferior 6-8 Transitional Area Area Frontal Opercular 5 $0.021$ HC 2.328
Frontal Opercular Area 1 Area Lateral IntraParietal dorsal $0.006$ HC 2.319
Area posterior 10p Area 7PC $0.021$ HC 2.310
Primary Auditory Cortex Area 7PC $0.023$ HC 2.308
Dorsal area 6 Area 7PC $0.021$ HC 2.305
Anterior Agranular Insula Complex Area Frontal Opercular 5 $0.023$ HC 2.290
ParaHippocampal Area 1 Area Lateral IntraParietal dorsal $0.007$ HC 2.287
Auditory 4 Complex Area IFJa $0.022$ HC 2.276
Auditory 4 Complex Area posterior 9-46 ventral $0.024$ HC 2.267
Area IFJa Area Frontal Opercular 5 $0.022$ HC 2.265
Area 9 Middle Area PGs $0.022$ HC 2.264
Area posterior 9-46 ventral Area Frontal Opercular 5 $0.024$ HC 2.256
Area anterior 9-46 ventral Area Frontal Opercular 5 $0.023$ HC 2.255
Area PF Opercular Area Frontal Opercular 5 $0.023$ HC 2.254
Frontal Opercular Area 1 Area 7PC $0.023$ HC 2.248
PreSubiculum Area PGs $0.023$ HC 2.247
Superior 6-8 Transitional Area Area Lateral IntraParietal dorsal $0.024$ HC 2.241
Area TA2 (temporal region A, area 2) Area Lateral IntraParietal dorsal $0.023$ HC 2.241
Area OP1-SII Area Frontal Opercular 5 $0.024$ HC 2.229
Area s32 Area 7PC $0.024$ HC 2.228
Medial Belt Complex Area Frontal Opercular 5 $0.008$ HC 2.226
Area PH Area IFJa $0.024$ HC 2.224
Area Posterior Insular 1 Area Lateral IntraParietal dorsal $0.024$ HC 2.215
VentroMedial Visual Area 1 Area Frontal Opercular 5 $0.025$ HC 2.213
Area 33 prime Area Frontal Opercular 5 $0.025$ HC 2.211
Area 43 Area IFJa $0.026$ HC 2.209
Area OP2-3-VS (Ventral-Superior part of 2-3 area of Parietal Operculum) Area Lateral IntraParietal dorsal $0.025$ HC 2.206
Frontal Opercular Area 1 Area Frontal Opercular 5 $0.025$ HC 2.205
Pirform Cortex Area Lateral IntraParietal dorsal $0.009$ HC 2.187
Area 9 anterior Area PGs $0.026$ HC 2.183
Area 8B Lateral Area PGs $0.026$ HC 2.183
Area PF Complex Area Frontal Opercular 5 $0.027$ HC 2.179
Anterior Agranular Insula Complex Area posterior 9-46 ventral $0.027$ HC 2.176
Area 1 Area Frontal Opercular 5 $0.027$ HC 2.173
Premotor Eye Field Area Frontal Opercular 5 $0.027$ HC 2.173
ParaHippocampal Area 1 Area Frontal Opercular 5 $0.027$ HC 2.172
Area PFt Area Frontal Opercular 5 $0.027$ HC 2.169
Area 13 lateral Area Frontal Opercular 5 $0.027$ HC 2.163
Rostral Area 6 Area 7PC $0.028$ HC 2.159
Anterior Agranular Insula Complex Area 7PC $0.029$ HC 2.143
Area 6 medial posterior Area PGs $0.028$ HC 2.139
Area 43 Area Frontal Opercular 5 $0.028$ HC 2.138
Area OP4-PV (Posterior-Ventral part of area 4 of Parietal Operculum) Area Lateral IntraParietal dorsal $0.028$ HC 2.138
Posterior Insular Area 2 Area posterior 9-46 ventral $0.029$ HC 2.128
ParaHippocampal Area 1 Sixth Visual Area $0.030$ HC 2.114
Parieto-Occipital Sulcus Area 1 Area Lateral IntraParietal dorsal $0.009$ HC 2.111
Area posterior 10p Area Frontal Opercular 5 $0.031$ HC 2.108
Ventral Area 6 Area Frontal Opercular 5 $0.030$ HC 2.106
Area STSd anterior Area Lateral IntraParietal dorsal $0.009$ HC 2.100
Area anterior 9-46 ventral Area Lateral IntraParietal dorsal $0.031$ HC 2.093
Area Posterior Insular 1 Area PGs $0.031$ HC 2.093
Rostral Area 6 Area Frontal Opercular 5 $0.010$ HC 2.091
Area FST Area Frontal Opercular 5 $0.031$ HC 2.085
Area Superior temporal sulcus ventral anterior Area 7PC $0.031$ HC 2.081
Area Posterior Insular 1 Area posterior 9-46 ventral $0.032$ HC 2.076
Area TE1 Middle Area 7PC $0.032$ HC 2.070
Area 6 medial posterior Area 7PC $0.032$ HC 2.069
Area Posterior Insular 1 Area 7PC $0.032$ HC 2.067
Area Posterior Insular 1 Area Frontal Opercular 5 $0.033$ HC 2.063
Area TA2 (temporal region A, area 2) Area PGs $0.032$ HC 2.062
PreSubiculum Area Lateral IntraParietal dorsal $0.011$ HC 2.055
Area 43 Area 7PC $0.033$ HC 2.048
ParaHippocampal Area 2 Area 7PC $0.034$ HC 2.045
Inferior 6-8 Transitional Area Area Lateral IntraParietal dorsal $0.034$ HC 2.044
ParaHippocampal Area 3 Area Lateral IntraParietal dorsal $0.012$ HC 2.043
Area 10 rostral Area PGs $0.034$ HC 2.041
ParaBelt Complex Area Lateral IntraParietal dorsal $0.034$ HC 2.041
Medial Superior Temporal Area Area Frontal Opercular 5 $0.034$ HC 2.041
Superior 6-8 Transitional Area Area PGs $0.034$ HC 2.040
Entorhinal Cortex Area Lateral IntraParietal dorsal $0.034$ HC 2.037
VentroMedial Visual Area 1 Area IFJa $0.034$ HC 2.036
Area Lateral Occipital 3 Area Lateral IntraParietal dorsal $0.012$ HC 2.033

### v2

@@ Appendix F Interpretability
Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
Area PHT L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 5.637
Area 9 Posterior R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 5.334
Area 8Ad R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.99
Area 8B Lateral L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.835
Orbital Frontal Complex L Area TF R ASD $<0.001$ 3.767
Superior Frontal Language Area L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.566
Superior 6-8 Transitional Area L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.566
Area TE2 posterior L Area Posterior 24 prime R ASD $<0.001$ 3.549
RetroInsular Cortex R Area anterior 32 prime L ASD $<0.001$ 3.54
Frontal Opercular Area 2 L Ventral IntraParietal Complex R ASD $<0.001$ 3.499
Fourth Visual Area L Area anterior 32 prime L ASD $<0.001$ 3.494
Lateral Area 7A R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.38
Frontal Eye Fields L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.355
Area V3B R Area 9-46d L ASD 0.002 3.339
Medial Area 7P L Area anterior 32 prime L ASD 0.002 3.328
Fourth Visual Area L Area PFm Complex L ASD 0.002 3.295
Sixth Visual Area R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.252
Area posterior 9-46v R Area anterior 32 prime L ASD $<0.001$ 3.201
Second Visual Area R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.176
Area 8BM R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.173
Area STSd posterior L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 3.169
Second Visual Area R Area 23d R ASD $<0.001$ 3.143
IntraParietal Sulcus Area 1 R Area 23d R ASD $<0.001$ 3.128
Lateral Area 7P L RetroInsular Cortex L ASD $<0.001$ 3.08
Lateral Area 7P R RetroInsular Cortex L ASD $<0.001$ 3.076
Medial Area 7P L Primary Sensory Cortex R ASD $<0.001$ 3.064
Second Visual Area L Area anterior 32 prime L ASD $<0.001$ 3.024
Second Visual Area L Area PFm Complex L ASD $<0.001$ 3.012
Area posterior 9-46v L RetroInsular Cortex L ASD $<0.001$ 3.009
Area TE2 posterior L Area 5m R ASD $<0.001$ 2.983
Area 5m L Area anterior 32 prime L ASD $<0.001$ 2.966
Area PH L Area anterior 32 prime L ASD $<0.001$ 2.963
Area 2 R Area anterior 32 prime L ASD $<0.001$ 2.962
Area V3B R Area anterior 32 prime L ASD $<0.001$ 2.95
Area 9 Middle L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.928
Area 31a R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.921
Perirhinal Ectorhinal Cortex R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.919
Area 6m anterior R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.913
Area PH L Area PFm Complex L ASD $<0.001$ 2.878
Ventral Area 6 R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.867
Parieto-Occipital Sulcus Area 1 L Medial Superior Temporal Area R ASD $<0.001$ 2.859
Ventral Visual Complex R Area anterior 32 prime L ASD $<0.001$ 2.859
Area anterior 32 prime R Area 23d R ASD $<0.001$ 2.844
Area TemporoParietoOccipital Junction 3 R Area anterior 32 prime L ASD $<0.001$ 2.838
ParaHippocampal Area 2 R Area anterior 32 prime L ASD $<0.001$ 2.821
Area 9-46d R Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.82
Area V3B R Area Posterior 24 prime R ASD $<0.001$ 2.801
Area PHT L Area dorsal 23 a+b L ASD $<0.001$ 2.801
Area 47m L Area Lateral Occipital 1 L ASD $<0.001$ 2.796
Ventral IntraParietal Complex L Medial Superior Temporal Area L ASD $<0.001$ 2.796
Area IntraParietal 2 R Area anterior 32 prime L ASD $<0.001$ 2.789
Area STSv posterior L Area 23d R ASD 0.007 2.781
VentroMedial Visual Area 1 L Primary Sensory Cortex R ASD $<0.001$ 2.775
Area posterior 24 L Area TemporoParietoOccipital Junction 3 L ASD $<0.001$ 2.774
Medial Area 7A R Primary Sensory Cortex R ASD $<0.001$ 2.767
Primary Motor Cortex R Area anterior 32 prime L ASD 0.001 2.762
Medial Area 7A L Primary Sensory Cortex R ASD 0.001 2.753
Medial Area 7P L Area PFm Complex L ASD $<0.001$ 2.751
Dorsal Transitional Visual Area L Medial Superior Temporal Area R ASD 0.001 2.744
Area Lateral Occipital 2 R RetroInsular Cortex L ASD 0.001 2.736
Area Lateral Occipital 2 R Medial Superior Temporal Area R ASD 0.001 2.734
Area dorsal 32 L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.723
Area TG Ventral L RetroInsular Cortex L ASD 0.001 2.711
Area Lateral IntraParietal dorsal R Area anterior 32 prime L ASD 0.001 2.696
Primary Visual Cortex R Area 23d R ASD 0.001 2.688
Area p32 R Medial Superior Temporal Area R ASD 0.005 2.685
Area STSv posterior L Area PFm Complex L ASD 0.005 2.683
Area PFt R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.681
Area 31a R Area PFm Complex L ASD 0.001 2.672
Area TemporoParietoOccipital Junction 3 R Area PFm Complex L ASD 0.001 2.669
Dorsal Transitional Visual Area R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.658
Frontal Opercular Area 3 L Area 8Av L ASD 0.001 2.645
Anterior IntraParietal Area R Area anterior 32 prime L ASD 0.001 2.627
Dorsal Transitional Visual Area R Area PFm Complex L ASD 0.001 2.62
Posterior Insular Area 2 R Area anterior 32 prime L ASD 0.001 2.618
Primary Visual Cortex L Area 23d R ASD 0.001 2.615
Medial Area 7A L Area anterior 32 prime L ASD 0.001 2.608
Primary Visual Cortex L Area PFm Complex L ASD 0.001 2.606
Area 8Av L Primary Sensory Cortex R ASD 0.001 2.599
Area TE2 posterior L RetroInsular Cortex L ASD 0.001 2.583
Area Lateral Occipital 2 R Area dorsal 23 a+b L ASD 0.001 2.58
RetroInsular Cortex R Area PFm Complex L ASD 0.001 2.578
Area 1 R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.576
Primary Sensory Cortex R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.572
Area STSv anterior L Area 6 anterior L ASD 0.001 2.57
Area V3B R Medial Belt Complex L ASD 0.001 2.568
Seventh Visual Area L Medial Superior Temporal Area L ASD 0.003 2.565
Area V3A R Ventral IntraParietal Complex R ASD 0.001 2.565
Area STGa L Area 23d R ASD 0.001 2.558
Area 2 L Ventral IntraParietal Complex R ASD 0.001 2.55
Para-Insular Area R Area 23d R ASD 0.001 2.548
Inferior 6-8 Transitional Area L RetroInsular Cortex L ASD 0.001 2.548
Seventh Visual Area L Area PFm Complex L ASD 0.001 2.546
Dorsal area 6 R Area 9-46d L ASD 0.001 2.541
Area 8BM L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.536
Area TE2 posterior L Area Posterior 24 prime L ASD 0.001 2.535
Area 52 L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.533
Area dorsal 32 R Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.532
Ventral Area 6 R Area 23d R ASD 0.001 2.53
Area 9 Posterior L Area TemporoParietoOccipital Junction 3 L ASD 0.001 2.53